# Time series forecasting of multiphase microstructure evolution using deep learning

### v1

@@ IV.4 Integration of surrogate and phase-field model frameworks
With proper training, the surrogate model takes a fraction of a second to forecast ten future frames, while the phase-field model takes about 20 seconds to simulate the same. Thus, integrating our data-driven approach with the phase-field solver makes an accelerated phase-field framework. In order to accelerate phase-field simulations, a given number of solver iterations can be replaced or “leaped in time” by the surrogate model predictions of corresponding timesteps. To demonstrate this application, we input the surrogate model predictions as an initial configuration to the phase-field solver to emulate microstructure evolution in the long-term. We note that the solver does not execute when the initial configuration is taken as frame $t+30$ and beyond, likely due to the considerable noise present in these frames that gets accumulated over time series forecasting. Thus, we can only simulate the evolution of the frames $t+10$ and $t+20$ in the phase-field solver (Fig. 10). We compare their long-term evolution with the true evolution from the full-scale phase-field simulation implemented with a random initial configuration. Comparing the evolution trajectory at $t=5000$, significant differences between the ground truth and predictions can be seen. However, these artifacts are eliminated in the long-term ($t=10000$), leading to excellent agreement between them.
The above observations demonstrate that we can effectively replace 20 timesteps from the phase-field solver by the surrogate model predictions and then, using the prediction (say, $t+10$), run the phase-field simulation for the next 20 timesteps, and so on. Such alternate use of phase-field and surrogate models to generate the full microstructure evolution process can lower the computing cost by at least 40 times relative to the full-scale phase-field simulation, even with no loss in accuracy. To further quantify the merits of this hybrid approach, we use a Fourier analysis that accurately captures the statistical properties of microstructures, including the domain size. We study the average domain size (a measure of characteristic length scale) of component $A$ in the microstructure by circularly-averaged structure functions [42, 43]:
@@
where $N$ is the lattice size, $c_{A}(\bm{k},t)$ is the Fourier transform of $c_{A}(\bm{r},t)$, and $k$ is the magnitude of the wave vector $\bm{k}$. The $S_{A}$ can capture statistically similar but quantitatively different local morphological features in $A$, with its maximum (say, $k_{\text{max}}$) determines the average domain size. As illustrated in Fig. 11, the domain size (given by $2\pi/k_{\text{max}}$) in frame $t=920$ (i.e., $t+10$) agrees well with the ground truth. However, this agreement is lost for future time series predictions (e.g., $t=1010$ and $t=1120$), likely due to the accumulated noise over time series forecasting, weakening the signal of the power spectrum (i.e., low peak height at $k_{\text{max}}$). Such quantitative analysis of spatial statistics confirms that we can input frame $t=920$ in the phase-field solver to produce the correct microstructure evolution trajectory without valuable information/accuracy loss on the salient physics of the phenomena of interest. We note that the MSE (Fig. 9) with the predicted frame $t=920$ is given by 0.002, which we can consider as benchmark MSE for the ternary system to characterize the quality of the forecasting accuracy. Such analyses further demonstrate the good predictive capability of our robust yet simple data-driven framework.
@@
IV.5 Transfer learning to forecast unseen composition and physics
Our surrogate model shows satisfactory performance in predicting the time series of microstructure evolution in the ternary mixture ($A_{1/4}B_{1/4}C_{1/2}$) described earlier (Sec. IV.2). Can we use this pre-trained surrogate model to forecast microstructures from different datasets/mixtures? This problem is equivalent to predicting future time series values of the unknown image sequence generated by different simulations.
We use two different test datasets unknown to the trained surrogate model. In the first dataset, we vary the material composition ($A_{1/3}B_{1/3}C_{1/3}$ in Fig. 12a1) that leads to new microstructure morphologies, specifically the red-colored phase now takes a globular shape compared to its continuous morphology in Fig. 2. However, in this case, the evolution principles remain the same: phase separation followed by the domain coarsening. In the second dataset, we vary material parameters ($\chi_{ij}$ and $\kappa_{i}$ in Eq. (2) following Ref. [31]) to introduce surface-energy-driven phenomena (e.g., wetting) in the phase-separating system in addition to the domain coarsening effects. The wetting case is simulated after setting the thermodynamic interactions between the mixture components asymmetrically as $\chi_{AB}$ = 2.5, $\chi_{BC}$ = 3.5, $\chi_{AC}$ = 5.0; and $\kappa_{A}$ = 2.0, $\kappa_{B}$ = 6.0, and $\kappa_{C}$ = 6.0 (Sec. II). The wetting here refers to the selective preference of forming the low-energy interface between coexisting phases. This is illustrated in Fig. 12b1, where wetting-induced composition layers of the preferred phase develop surrounding the circular particle phase. Such “ring pattern” formation during spinodal decomposition was observed in experiments [44, 45].
@@ IV.5 Transfer learning to forecast unseen composition and physics
Figure 12: (Color online) Referring to Sec. IV.5, we use our pre-trained surrogate model for forecasting different time series of microstructure evolution produced with two different scenarios in mind: (a1) different material compositions without altering the spinodal decomposition physics known to the surrogate model, and (b1) altering parameters that lead to new physical mechanisms (e.g., wetting phenomenon) unknown to the model. We use the one-frame ahead forecasting (similar to Fig. 5) as a baseline reference to investigate the performance of the surrogate model on unseen microstructures generated from different simulations. (a2) On average, the MSE remains low for predictions on unseen composition with no new physics involved. (b2) The MSE becomes high when new physics is present in the system for which evolution takes a completely different pathway from the model’s knowledge. Typically, the early and late time trends in the MSE curves correspond to the rapid and slow regimes of microstructure evolution. We use colors for visualization of binary (blue: $A$ and yellow: $B$) and ternary (green: $A$, blue: $B$, and red: $C$) microstructures.
V Conclusion and perspectives
@@
  • We have shown that RNN models can correctly predict the short-term spatiotemporal dynamics of microstructure evolution, but their forecasting ability decreases in the long-term. Also, the MSE calculated between ground truth and predictions with the RNN models increasingly deviate from each other over time. Among all RNN models, LSTM demonstrates the best learning ability, prediction accuracy, and long-term forecasting ability with the lowest MSE.
@@
  • Despite the satisfactory performance of all RNN models (i.e., low MSE), their learning rate and predictive power vary with the nature of the microstructure evolution problem to which they are applied. The agreement between the predicted and true microstructures is better in the ternary system than in the binary system. These observations can be attributed to the time evolution of the microstructure spatial statistics with drastic topological changes, including inherent randomness during binary phase separation, and the role of the “third” element in the ternary system that provides a barrier to long-range diffusional transport processes, inducing a local ordering in the microstructure that leads to a more slow, steadier microstructure evolution.
@@
  • With proper training, our surrogate model makes time series forecasting in a fraction of a second with increasing loss in accuracy over time. Using temporal extrapolation tests for multi-frame ahead forecasting (Fig. 8) and the Fourier analysis of the predictions (Fig. 11), we demonstrate how many time series values can be used as an initial configuration to the phase-field solver for generating correct microstructure evolution trajectory. Consequently, these predictions can effectively replace the corresponding timesteps in the phase-field solver, which takes significant time to compute. In this way, the phase-field and surrogate models can be integrated to significantly accelerate the microstructure evolution workflow with negligible loss in accuracy.
@@
  • We have tested the general ability of our surrogate model to forecast new microstructure evolution processes unknown to the model (i.e., to predict future values of different image sequences). We use an approach similar to transfer learning in which the surrogate model is trained with a known ternary dataset (Fig. 2) and subsequently applied to predict new microstructures created with different material compositions or underlying evolution physics. The RNN models, particularly LSTM, show quantitative prediction accuracy (MSE < 0.005) on new microstructures with unknown composition (Fig. 12). However, the agreement is qualitative (MSE < 0.01) when the model forecasts evolution on new microstructures with unknown physics. Qualitatively speaking, these examples imply our model’s intuition into the underlying physics and alloy chemistry of the microstructure formation, making it a useful general-purpose alternative to physics-based simulations of microstructure evolution.
Typical deep learning models do not incorporate physical rules of the spatiotemporal process. Also, high-fidelity phase-field data are expensive or time-consuming to obtain. Physics-informed deep models such as physics-informed neural networks [50] and physics-informed Gaussian processes [51] address both of these problems by enforcing physical constraint terms from known physical laws (e.g., continuity, conservation of mass) to an objective function. Moreover, deep learning models are essentially “black box” models with no guidance on the measurement uncertainty and on the input parameters essential for forecasting. Thus, a fusion of deep models with classical statistical models such as variational Bayesian inference, Monte Carlo dropout, and mixture density networks could be beneficial [14, 52]. Another common issue with the deep models, interpretability, can be improved, for example, using non-negative matrix factorization that can decompose the microstructure data into additive parts, which might correspond to meaningful physical components of the microstructure [53, 54].
Further, the readily parallelizable nature of convolutional operation makes convolutional CNNs and RNNs suitable for implementation on graphics processing unit (GPU). Thus, taking advantage of deep learning methods powered by GPU computing can massively accelerate the model training and forecasting, leading to rapid mapping of parameter-microstructure relations based on the time series prediction [57].
Phase-field simulations are used for quantitative microstructure evolution that accurately mimic experimental observations. Thus, phase-field datasets are considered as high-fidelity training data for developing predictive deep learning frameworks for forecasting a variety of microstructure evolution processes in diverse material systems. With this motivation, we develop the phase-field-informed surrogate model reported in this study that can be applied more generally to predict rich microstructure evolution problems due to spinodal decomposition, grain growth, order-disorder transformation, and alloy solidification, among others [29]. Undoubtedly, including experimental measurements within the training data will further enrich the classes of microstructure evolution. Moreover, we can use our framework to study microstructure statistics requiring a design-of-experiments approach [58, 59] in which the model and material parameters vary within predefined ranges to address the increasing data demand to predict microstructure evolution problems for unknown values of the input configuration such as alloy composition not directly “seen” by the model. It would also be interesting to use our framework to predict microstructure evolution in 3D since the diffusional processes and resultant length scale of the microstructures are very different in 3D compared to 2D even when all other parameters remain the same [11]. Finally, our framework will be increasingly improved based on the above recommendations, enabling it to handle more complexity in material microstructures with stronger spatiotemporal correlations between them along their evolution trajectory.
Acknowledgements.

### v2

@@ 4.4 Integration of surrogate and phase-field model frameworks
With proper training, the surrogate model takes a fraction of a second to forecast ten future frames, while the phase-field model takes about 20 seconds to simulate the same. Thus, integrating our data-driven approach with the phase-field solver makes an accelerated phase-field framework. In order to accelerate phase-field simulations, a given number of solver iterations can be replaced or “leaped in time” by the surrogate model predictions of corresponding timesteps. However, the question remains on how many phase-field steps can be replaced by the surrogate model to reproduce quantitative microstructure evolution trajectory without loss in accuracy. To demonstrate this approach, we input the surrogate model predictions as an initial configuration to the phase-field solver to emulate microstructure evolution in the long-term. We note that the solver does not execute when the initial configuration is taken as frame $t+30$ and beyond, likely due to the considerable noise present in these frames that gets accumulated over time series forecasting. Thus, we can only simulate the evolution of the frames $t+10$ and $t+20$ in the phase-field solver (Fig. 11). We compare their long-term evolution with the true evolution from the full-scale phase-field simulation implemented with a random initial configuration. Comparing the evolution trajectory at $t=5000$, significant differences between the ground truth and predictions can be seen. However, these artifacts are eliminated in the long-term ($t=10000$), leading to excellent agreement between them.
@@
The above observations demonstrate that we can effectively replace 20 timesteps from the phase-field solver by the surrogate model predictions and then, using the prediction (say, $t+10$), run the phase-field simulation for the next 20 timesteps, and so on. Such alternate use of phase-field and surrogate models to generate the full microstructure evolution process can lower the computing cost by at least 40 times relative to the full-scale phase-field simulation, even with no loss in accuracy. To further quantify the merits of this hybrid approach, we use the structure factor analysis (Eq. (4)) that accurately captures the statistical properties of microstructures, including the domain size. The $S_{A}$ in Eq. (4) can capture statistically similar but quantitatively different local morphological features in $A$, with its maximum (say, $k_{\text{max}}$) signifies a measure of the average domain size. As illustrated in Fig. 12, the domain size (given by $2\pi/k_{\text{max}}$) in frame $t=920$ (i.e., $t+10$) agrees well with the ground truth. However, this agreement is lost for future time series predictions (e.g., $t=1010$ and $t=1120$), likely due to the accumulated noise over time series forecasting, weakening the signal of the power spectrum (i.e., low peak height at $k_{\text{max}}$). Such quantitative analysis of spatial statistics confirms that we can input frame $t=920$ in the phase-field solver to produce the correct microstructure evolution trajectory without valuable information/accuracy loss on the salient physics of the phenomena of interest. We note that the MSE (Fig. 10) with the predicted frame $t=920$ is given by 0.002, which we can consider as benchmark MSE for the ternary system to characterize the quality of the forecasting accuracy. Such analyses further demonstrate the good predictive capability of our robust yet simple data-driven framework.
@@
4.5 Transfer learning to forecast unseen composition and physics
Our surrogate model shows satisfactory performance in predicting the time series of microstructure evolution in the ternary mixture ($A_{1/4}B_{1/4}C_{1/2}$) described earlier (Sec. 4.2). However, for a comprehensive understanding of the model behavior, it would be interesting to evaluate the predictive performance of the model on new microstructures generated using different material compositions or under different physical mechanisms. Thus, the question arises whether can we use the properly-trained surrogate model to forecast microstructures from different datasets/mixtures? This problem is equivalent to predicting future time series values of the unknown image sequence generated by different simulations.
We use two different test datasets unknown to the trained surrogate model. In the first dataset, we vary the material composition ($A_{1/3}B_{1/3}C_{1/3}$ in Fig. 13a1) that leads to new microstructure morphologies, specifically the red-colored phase now takes a globular shape compared to its continuous morphology in Fig. 2. However, in this case, the evolution principles remain the same: phase separation followed by the domain coarsening. In the second dataset, we vary material parameters ($\chi_{ij}$ and $\kappa_{i}$ in Eq. (2) following Ref. [31]) to introduce surface-energy-driven phenomena (e.g., wetting) in the phase-separating system in addition to the domain coarsening effects. The wetting case is simulated after setting the thermodynamic interactions between the mixture components asymmetrically as $\chi_{AB}$ = 2.5, $\chi_{BC}$ = 3.5, $\chi_{AC}$ = 5.0; and $\kappa_{A}$ = 2.0, $\kappa_{B}$ = 6.0, and $\kappa_{C}$ = 6.0 (Sec. 2). The wetting here refers to the selective preference of forming the low-energy interface between coexisting phases. This is illustrated in Fig. 13b1, where wetting-induced composition layers of the preferred phase develop surrounding the circular particle phase. Such “ring pattern” formation during spinodal decomposition was observed in experiments [44, 45].
Figure 13: (Color online) Referring to Sec. 4.5, we use our pre-trained surrogate model for forecasting different time series of microstructure evolution produced with two different scenarios in mind: (a1) different material compositions without altering the spinodal decomposition physics known to the surrogate model, and (b1) altering parameters that lead to new physical mechanisms (e.g., wetting phenomenon) unknown to the model. We use the one-frame ahead forecasting (similar to Fig. 6) as a baseline reference to investigate the performance of the surrogate model on unseen microstructures generated from different simulations. (a2) On average, the MSE remains low for predictions on unseen composition with no new physics involved. (b2) The MSE becomes high when new physics is present in the system for which evolution takes a completely different pathway from the model’s knowledge. Typically, the early and late time trends in the MSE curves correspond to the rapid and slow regimes of microstructure evolution. We use colors for visualization of binary (blue: $A$ and yellow: $B$) and ternary (green: $A$, blue: $B$, and red: $C$) microstructures.
@@ 4.5 Transfer learning to forecast unseen composition and physics
5 Discussion
Phase-field simulations are used for quantitative microstructure evolution that accurately mimic experimental observations. Thus, phase-field datasets are considered as high-fidelity training data for developing predictive deep learning frameworks for forecasting a variety of microstructure evolution processes in diverse material systems. With this motivation, we develop the phase-field-informed surrogate model reported in this study that can be applied more generally to predict rich microstructure evolution problems due to spinodal decomposition, grain growth, order-disorder transformation, and alloy solidification, among others [29]. Undoubtedly, including experimental measurements within the training data will further enrich the classes of microstructure evolution.
As mentioned in Sec. 1, spatiotemporal prediction of microstructure evolution using convolution RNNs has been studied considerably in the literature albeit for binary systems [16, 17, 18]. Also, the academic knowledge regarding the variance in efficacy when using different RNN variants in the surrogate model to predict the high-fidelity phase-field simulation is limited. It is well-established by these studies and others [20, 19, 22, 21] that several phase-field simulation timesteps can be effectively replaced by deep model predictions in an alternating hybrid approach between these models, considerably accelerating the phase-field simulation framework. However, these studies do not demonstrate how precisely the predicted frames can be applied to generate microstructure evolution in the long term. Moreover, for a more complete picture, can we predict the evolution in a new dataset of “unseen” microstructures with new physics unknown to the trained model? To our knowledge, these questions have not been addressed by previous studies. Moreover, there has been no study on the applications of deep models for multiphase systems.
As a case study, we use multi-phase-field model-generated datasets of spinodally decomposing ternary microstructures for data analysis. Additionally, we present the typical academic example of binary spinodal commonly reported in the literature as a reference. The present study addresses three novel questions: How does the surrogate model predict the future frames in two-phase and three-phase spinodal systems? How can we quantitatively (e.g., through power spectrum analysis) determine the phase-field simulation steps that can be “leaped in time” by deep model predictions for obtaining accurate long-term evolution? How accurately does our surrogate model predict the evolution history for unseen ternary microstructures simulated using new alloy composition and, particularly, physics?
We must emphasize that we have used different extensions of the free energy functions when comparing the binary (Eq. (1)) and ternary systems (Eq. (2)), which may lead to qualitative differences. Unfortunately, the present ternary formulation can not be applied to simulate binary systems due to the constraints that need to be satisfied for which complete freezing of one of the components is not possible (see [31] for details). However, we can approximate the phase-field model to work with a critical ($50:50$) binary composition but with a negligible amount of the third component; such preliminary analysis (Fig. 14) further confirms the general trends of lower MSE in predicting ternary systems, similar to the observations reported in Fig. 7.
Figure 14: We compare the prediction accuracy of the surrogate model between the ternary system and two different binary systems (or free energy functions). The binary system-1 is simulated following Eq. (1), and binary system-2 is implemented by approximating the ternary formulation in Eq. (2) to simulate a critical binary composition. For reference, we use LSTM for the next frame forecasting following Sec. 4.2.
We acknowledge that the current implementation, while using a simple semi-implicit solver, serves as a reasonable first step for coupling surrogate models with traditional phase-field simulations. The focus here is on demonstrating how a surrogate model can effectively accelerate phase-field simulations, even with a basic solver utilizing a small time step. We clarify that the acceleration from surrogate models comes not just from a single-step improvement but from eliminating the need to compute multiple time steps explicitly in a phase-field solver, reducing the overall simulation time. The RNN learns the dynamics and predicts future states directly, bypassing many intermediate time steps. It is straightforward to understand that this acceleration effect will be manifold when working with large-scale systems to obtain long-term evolution history. Future work will explore combining the surrogate model with implicit or more advanced methods [5], which could allow for larger time step size, reducing the number of simulation time steps needed without sacrificing numerical accuracy. This gradual improvement will further optimize the phase-field solver and surrogate model for speed and efficiency.
We must admit that we accelerated the numerical simulations efficiently using the available computing resources. For example, we simulated the phase-field simulations in parallel utilizing eight threads in an AMD Reyzen 5600 CPU operating at 4.4 GHz supported by 16 GB of RAM. We then moved the generated data to the GPU-accelerated Kaggle [39] environment (NVIDIA T4 $\times$ 2) for data-driven analysis. It is to be noted that the readily parallelizable nature of convolutional operation makes convolutional CNNs and RNNs more suitable for implementation on GPUs [46]. We emphasize that our surrogate model needs to be trained once before it can be used to predict future frames without further retraining. On average, the training time was $\approx$124 s, while the inference time for predicting a future frame using the trained model was $\approx$0.8 s. Undoubtedly, when more CPU/GPU resources become available, these times will be further reduced. Moreover, for the given ternary spinodal problem, we have found that surrogate predictions can safely replace 20 phase-field steps without sacrificing the prediction accuracy regarding the long-term microstructure evolution. For a simulation domain size of $512\times 512$, the phase-field calculation typically takes 40 s to solve 20 timesteps, while the surrogate model takes $\approx$0.8 s to predict the same number of future frames, leading to a speedup ratio of 50 (excluding the training time). This gain in computing time will accumulate when alternate phase-field and surrogate solvers will be used to expedite the entire simulation workflow. With increasing domain size, the speedup in computing time will become even more evident. We acknowledge that, among others, using more advanced numerical solution methods in the phase-field simulations and the type and number of threads in parallel computing framework will affect the computing cost and, hence, the speedup ratio. We will contemplate these problems in the future.
Typical deep models do not incorporate physical rules (e.g., partial differential equations) of the spatiotemporal process. Thus, RNNs can capture “new” microstructure patterns, possibly related to unknown physics, purely through data-driven learning. Although RNNs are not explicitly informed by governing equations, they can learn complex patterns and temporal dependencies from large datasets, effectively approximating unknown physical phenomena, which may or may not fully align with true underlying physics. In this context, physics-informed deep models such as physics-informed neural networks [51] and physics-informed Gaussian processes [52] explicitly incorporate known governing equations (e.g., continuity, conservation of mass) into the training process, ensuring that the predictions follow the fundamental physical laws. Moreover, deep learning models are essentially “black box” models with no guidance on the measurement uncertainty and on the input parameters essential for forecasting. Thus, a fusion of deep models with classical statistical models such as variational Bayesian inference, Monte Carlo dropout, and mixture density networks could be beneficial [14, 53]. Another common issue with the deep models, interpretability, can be improved, for example, using non-negative matrix factorization that can decompose the microstructure data into additive parts, which might correspond to meaningful physical components of the microstructure [54, 55].
Moreover, we can use our framework to study microstructure statistics requiring a design-of-experiments approach [58, 59] in which the model and material parameters vary within predefined ranges to address the increasing data demand to predict microstructure evolution problems for unknown values of the input configuration such as alloy composition not directly “seen” by the model. It would also be interesting to use our framework to predict microstructure evolution in 3D since the diffusional processes and resultant length scale of the microstructures are very different in 3D compared to 2D even when all other parameters remain the same [11]. Finally, our framework will be increasingly improved based on the above recommendations, enabling it to handle more complexity in material microstructures with stronger spatiotemporal correlations between them along their evolution trajectory.
6 Conclusions and outlook
@@
  • We have shown that RNN models can correctly predict the short-term spatiotemporal dynamics of microstructure evolution, but their forecasting ability decreases in the long-term. Also, the MSE calculated between ground truth and predictions with the RNN models increasingly deviates from each other over time. Among all RNN models, LSTM demonstrates the best learning ability, prediction accuracy, and long-term forecasting ability with the lowest MSE due to its ability to retain long-term dependencies in data due to the specialized gating mechanisms (input, forget, and output gates) to manage the flow of information. Other RNN variants, such as simple RNN and GRU, are expected to perform similarly when modeling short sequences, small datasets, or data with minimal noise, where long-term dependencies are not crucial. However, these effects have not been tested presently in this study.
@@
  • Despite the satisfactory performance of all RNN models (i.e., low MSE), their learning rate and predictive power vary with the nature of the microstructure evolution problem to which they are applied. The agreement between the predicted and true microstructures is better in the ternary system than in the binary system. These observations can be attributed to the time evolution of the microstructure spatial statistics with drastic topological changes, including inherent randomness during binary phase separation, and the role of the “third” element in the ternary system that may provide a barrier to long-range diffusional transport processes, leading to a more slow, steadier microstructure evolution compared to the binary systems (Fig. 5).
@@
  • With proper training, our surrogate model makes time series forecasting in a fraction of a second with increasing loss in accuracy over time. Using temporal extrapolation tests for multi-frame ahead forecasting (Fig. 9) and the Fourier analysis of the predictions (Fig. 12), we demonstrate how many time series values can be used as an initial configuration to the phase-field solver for generating correct microstructure evolution trajectory. Consequently, these predictions can effectively replace the corresponding timesteps in the phase-field solver, which takes significant time to compute. In this way, the surrogate-model assisted phase-field sovler can significantly accelerate the microstructure evolution workflow with negligible loss in accuracy.
@@
  • We have tested the general ability of our surrogate model to forecast new microstructure evolution processes unknown to the model (i.e., to predict future values of different image sequences). We use an approach similar to transfer learning in which the surrogate model is trained with a known ternary dataset (Fig. 2) and subsequently applied to predict new microstructures created with different material compositions or underlying evolution physics. The RNN models, particularly LSTM, show quantitative prediction accuracy (MSE < 0.005) on new microstructures with unknown composition (Fig. 13). However, the agreement is qualitative (MSE < 0.01) when the model forecasts evolution on new microstructures with unknown physics. Qualitatively speaking, these examples imply our model’s intuition into the underlying physics and alloy chemistry of the microstructure formation, making it a useful general-purpose alternative to physics-based, PDE-based approaches to microstructure evolution.
Acknowledgments